# UniSim: A Unified Simulator for Time-Coarsened Dynamics of Biomolecules

### v1

@@ 3.2 Unified Pretraining
${\mathcal{G}}_{g}$ =$=\{j|j\in{\mathcal{G}},||\vec{{\bm{x}}}_{j}-\vec{{\bm{x}}}_{c}||_{2}<\delta_{\mathrm{min}}\},$ (2)
${\mathcal{G}}_{e}$ =$=\{j|j\in{\mathcal{G}},||\vec{{\bm{x}}}_{j}-\vec{{\bm{x}}}_{c}||_{2}<\delta_{\mathrm{max}}\}.$ (3)
@@
${\mathcal{L}}_{e}=||\nabla_{\vec{{\bm{X}}}}(\sum_{i}{\bm{H}}^{\prime}_{\mathrm{out}}[i])-\frac{\vec{{\bm{X}}}^{\prime}-\vec{{\bm{X}}}}{\sigma_{e}}||_{2}^{2}.$ (10)
@@ 3.3 Vector Field Model for Dynamics
${\mathcal{L}}_{v}=\mathbb{E}[||v(t,\vec{{\mathbf{X}}}_{t})-(\vec{{\mathbf{X}}}_{1}-\vec{{\mathbf{X}}}_{0})||_{2}^{2}+||\eta_{z}(t,\vec{{\mathbf{X}}}_{t})-\vec{{\mathbf{Z}}}||_{2}^{2}],$ (17)
@@ 3.4 Force Guidance Kernel for Finetuning
where $g(\vec{{\mathbf{X}}}_{0},\vec{{\mathbf{X}}}_{1})=\exp(-\alpha(\varepsilon(\vec{{\mathbf{X}}}_{0})+\varepsilon(\vec{{\mathbf{X}}}_{1})))$ and $h(\vec{{\mathbf{X}}}_{0},\vec{{\mathbf{X}}}_{1},\vec{{\mathbf{X}}}_{t})=\nabla\log{q_{t}(\vec{{\mathbf{X}}}_{t})}-\nabla\log{q_{t}(\vec{{\mathbf{X}}}_{t}|\vec{{\mathbf{X}}}_{0},\vec{{\mathbf{X}}}_{1})}$. According to Albergo et al. (2023), $-\gamma^{-1}(t)\eta_{z}(t,\vec{{\mathbf{X}}}_{t})$ is an unbiased estimation of $\nabla\log{q_{t}(\vec{{\mathbf{X}}}_{t})}$, and $\nabla\log{q_{t}(\vec{{\mathbf{X}}}_{t}|\vec{{\mathbf{X}}}_{0},\vec{{\mathbf{X}}}_{1})}$ has a closed-form solution based on Eqs. 12 and 16, thus all terms of Eq. 20 can be calculated during training without any approximation.
@@ Appendix B Proofs of Propositions
According to the above derivation, by replacing $\nabla\log{p_{t}(\vec{{\mathbf{X}}})},\nabla\log{q_{t}(\vec{{\mathbf{X}}})}$ with $-\gamma^{-1}(t)\eta_{z}^{\prime}(t,\vec{{\mathbf{X}}}),-\gamma^{-1}(t)\eta_{z}(t,\vec{{\mathbf{X}}})$, respectively, we show that $\eta_{z}^{\prime}(t,\vec{{\mathbf{X}}})=\eta_{z}(t,\vec{{\mathbf{X}}})+\alpha\gamma(t)\nabla\varepsilon_{t}(\vec{{\mathbf{X}}})$ holds.
@@ C.1 Normalization of Potential and Force Labels
where $\max(\cdot)$ and $\min(\cdot)$ denote the maximum and minimum element of the vector. Based on the technique, all potential energy labels are normalized to $[-1,0]$ without manipulating the energy distribution of each ensemble.
@@ D.1 Dataset Details
Frame spacing 0.5fs 1ps 10ps
Simulation time per traj not fixed 100ns 10ns
Coarsened time $\tau$ 100ps 10ps 100ps

### v3

@@ 3.2 Unified Pretraining
[…]
@@
[…]
@@ 3.3 Vector Field Model for Dynamics
[…]
@@ 3.4 Force Guidance Kernel for Finetuning
where $g(\vec{{\mathbf{X}}}_{0},\vec{{\mathbf{X}}}_{1})=\exp(-\alpha(\varepsilon(\vec{{\mathbf{X}}}_{0})+\varepsilon(\vec{{\mathbf{X}}}_{1})))$ and $h(\vec{{\mathbf{X}}}_{0},\vec{{\mathbf{X}}}_{1},\vec{{\mathbf{X}}}_{t})=\nabla\log{q_{t}(\vec{{\mathbf{X}}}_{t})}-\nabla\log{q_{t}(\vec{{\mathbf{X}}}_{t}|\vec{{\mathbf{X}}}_{0},\vec{{\mathbf{X}}}_{1})}$. According to Albergo et al. (2023), $\gamma^{-1}(t)\eta_{z}(t,\vec{{\mathbf{X}}}_{t})$ is an unbiased estimation of $\nabla\log{q_{t}(\vec{{\mathbf{X}}}_{t})}$, and $\nabla\log{q_{t}(\vec{{\mathbf{X}}}_{t}|\vec{{\mathbf{X}}}_{0},\vec{{\mathbf{X}}}_{1})}$ has a closed-form solution based on Eqs. 12 and 16, thus all terms of Eq. 20 can be calculated during training without any approximation.
@@ Appendix B Proofs of Propositions
According to the above derivation, by replacing $\nabla\log{p_{t}(\vec{{\mathbf{X}}})},\nabla\log{q_{t}(\vec{{\mathbf{X}}})}$ with $\gamma^{-1}(t)\eta_{z}^{\prime}(t,\vec{{\mathbf{X}}}),-\gamma^{-1}(t)\eta_{z}(t,\vec{{\mathbf{X}}})$, respectively, we show that $\eta_{z}^{\prime}(t,\vec{{\mathbf{X}}})=\eta_{z}(t,\vec{{\mathbf{X}}})+\alpha\gamma(t)\nabla\varepsilon_{t}(\vec{{\mathbf{X}}})$ holds.
@@ C.1 Normalization of Potential and Force Labels
where $\max(\cdot)$ and $\min(\cdot)$ denote the maximum and minimum element of the vector. Based on the technique, all potential energy labels are normalized to [1,0][-1,0] without manipulating the energy distribution of each ensemble.
@@ D.1 Dataset Details
Frame spacing 0.5fs 1ps 100ps
Simulation time per traj not fixed 100ns 100ns
Coarsened time $\tau$ 100ps 10ps 1ns